OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4. Answers: y^2 = -8x 16y = x^2 y = -1/16 x^2 x = -1/16 y^2

OpenStudy (anonymous):

@ant501 help on this one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Thanks lol another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@ant501 Find the vertex, focus, directrix, and focal width of the parabola. x = 3y^2 The answers are: vertex:(0,0) focus (1/12, 0) directrix: x = 1/12 focal width: 12 vertex (0,0) focus (1/12, 0) directrix: x = -1/12 focal with: 0.33 vertex (0,0) focus (0, 1/16) directrix x = -1/16 focal width: 0.33 vertex (0,0) focus (1/16, 0) directrix y = -1/16 focal width: 12

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i mean b

OpenStudy (anonymous):

That's the one I thought it was. One final one? I'll post a new question so I can give you more medals for the help. @ant501

OpenStudy (anonymous):

sure